the lengths of DG, EF and FG are given in units. The is the area, in square units of Triangle DEG?
DE=12
EF=7
FG=10

Answer by unlockmath(1688)

Hello,
We can find the area of triangle DEF by using the hypotenuse Formula.
DF(squared) + EF(squared)= DE(squared)
Plug in the numbers.
DF(squared) + 7(squared) = 12(squared)
Do the math.
DF(sq)+ 49 = 144
Subtract 49 from both sides.
DF (sq) = 95
Square root both sides.
DF = sq root of 95
Now plug in the area formula for that triangle which is 1/2 Base times height
1/2(7)sq rt 95 or

3.5 sq rt 95 we can approx sq rt of 95 to 9.75
Therefore: area of this triangle is 34.125 Sq Units
Now plug in the area formula for triangle DFG.
1/2(9.75)10 = Area
Do the math.
48.75 Square Units
Add both triangles and you have:
82.875 square units
RJ Toftness
Author "Unlock the Mystery to Math"
www.math-unlock.com